# HOT3D: Hand and Object Tracking in 3D from Egocentric Multi-View Videos

### v1

@@ Abstract
We introduce HOT3D, a publicly available dataset for egocentric hand and object tracking in 3D. The dataset offers over 833 minutes (more than 3.7M images) of multi-view RGB/monochrome image streams showing 19 subjects interacting with 33 diverse rigid objects, multi-modal signals such as eye gaze or scene point clouds, as well as comprehensive ground-truth annotations including 3D poses of objects, hands, and cameras, and 3D models of hands and objects. In addition to simple pick-up/observe/put-down actions, HOT3D contains scenarios resembling typical actions in a kitchen, office, and living room environment. The dataset is recorded by two head-mounted devices from Meta: Project Aria, a research prototype of light-weight AR/AI glasses, and Quest 3, a production VR headset sold in millions of units. Ground-truth poses were obtained by a professional motion-capture system using small optical markers attached to hands and objects. Hand annotations are provided in the UmeTrack and MANO formats and objects are represented by 3D meshes with PBR materials obtained by an in-house scanner. In our experiments, we demonstrate the effectiveness of multi-view egocentric data for three popular tasks: 3D hand tracking, 6DoF object pose estimation, and 3D lifting of unknown in-hand objects. The evaluated multi-view methods, whose benchmarking is uniquely enabled by HOT3D, significantly outperform their single-view counterparts.
@@ 1 Introduction
Figure 2: Sample images from Aria (top) and Quest 3 (bottom). Aria recordings include one RGB and two monochrome image streams, while Quest 3 recordings include two monochrome streams – only images from one of the multi-view streams are shown. Contours of 3D models of hands and objects in the ground-truth poses are shown in white and green respectively. In addition to simple pick-up/observe/put-down actions, the subjects perform actions that are common in a kitchen, office, and living room. To increase diversity, the lighting, furniture, and decorations in the capture lab were regularly randomized.
@@
To accelerate computer vision research on hand-object interaction, we are publicly releasing HOT3D, an egocentric dataset for training and evaluating methods for hand and object tracking in 3D. The dataset is recorded using two recent head-mounted devices from Meta: Project Aria [13], which is a research prototype of light-weight AI glasses, and Quest 3 [41], a virtual-reality headset that has shipped millions of units. The dataset offers over 833 minutes of egocentric video streams, which include over 1.5M multi-view frames (over 3.7M images) and show 19 subjects interacting with 33 diverse rigid objects. Besides a simple inspection scenario, where subjects pick up, observe, and put down the objects, the sequences show scenarios resembling typical actions in kitchen, office, and living room spaces. Hands and objects are annotated with accurate 3D poses collected using a passive marker-based motion-capture system. The dataset also includes 3D object models which were obtained by an in-house scanning-based 3D object reconstruction pipeline and include high-resolution geometry and PBR materials [40]. Recordings from Aria additionally include 3D scene point clouds from SLAM and eye gaze information. Sample images are in Fig. 2.
The HOT3D dataset is primarily intended for training and evaluation of hand and object tracking methods in 3D space from localized, egocentric, multi-view video streams, as opposed to monocular views or individual images. Since images from all streams are synchronized with a hardware trigger (i.e., captured at the same timestamp), the dataset enables development of methods that can leverage multi-view and/or temporal information. While providing a testbed for CAD-based object tracking, the dataset also enables a CAD-free tracking setup by providing reference sequences showing different views at each object, which can be used to onboard the objects in a few-shot manner. Furthermore, the dataset can be used for tasks such as 3D object reconstruction and 2D detection or segmentation of hand-object interactions. We also encourage research that leverages the eye gaze information from Aria, which can enable predicting the user’s intent, or efficient allocation of the computational budget via foveated sensing.
@@
ENIGMA-51 [52] 45K RGB 1 / 0 HoloLens 2 19 Both 25 25 Manual
@@
Table 1: Datasets with 3D hands and object annotations. HOT3D is the first dataset to provide multi-view, hardware time-synced egocentric videos captured with real headsets. It is the largest dataset in terms of total image count and provides high quality ground-truth annotations (comparable to [15]).
To analyze the effectiveness of multi-view egocentric data, we present experiments showing that multi-view methods significantly outperform their single-view variants. In addition to multi-view 3D object tracking, we focused on the tasks of multi-view 6DoF pose estimation of unseen objects and 3D lifting of in-hand objects [51], for which we developed strong baselines. Our results highlight the advantages of multi-view and temporal data, which is under-explored in hand/object tracking research due to limited datasets.
@@
    Strong baselines for tasks enabled by HOT3D: We developed simple yet powerful multi-view baselines for two increasingly popular tasks that are relevant for augmented/mixed reality: 6DoF pose estimation of objects with known 3D models (extending FoundPose [50]), and 3D lifting of unknown in-hand objects (based on stereo matching of DINOv2 [49] features). Source code of the methods will be publicly available.
@@
    Demonstrated effectiveness of multi-view egocentric data: Our experiments show that multi-view methods for 3D hand tracking, 6DoF object pose estimation, and 3D lifting of handheld objects clearly outperform their single-view counterparts. This is an important result for research on power-efficient egocentric vision systems, which can typically afford a multi-view camera setup [13] but not e.g., active depth sensors [68].
@@ 2 Related work
The progress of research in computer vision has been strongly influenced by benchmark datasets [58, 14, 57, 19, 32] which enable comparing of methods and understanding of their limitations. In this section, we first review existing datasets with either hand or object pose annotations, and then focus on datasets that offer annotations of hands and hand-manipulated objects.
Datasets with hands only. Vision-based 3D hand pose estimation and tracking has been extensively studied for many years, with the first methods focusing on custom datasets with monochrome images [54, 29]. Significant improvements in pose accuracy were later achieved on RGB-D images from datasets such as NYU [66], ICVL [65], MSRA [63], Tzionas et al. [67], EgoDexter [44], or HANDS17 [73]. Recently, partly motivated by AR/VR use cases where depth sensors are often unavailable due to high power consumption, the research community has largely switched to RGB or monochrome images, working on datasets such as the Stereo dataset [74], InterHand2.6M [43], FreiHAND [76], UmeTrack [26], AssemblyHands [47], and datasets with pose annotations of both hands and objects reviewed below.
@@
Datasets with hands and objects. Many existing datasets include images of hands and objects (e.g., [16, 6, 1, 60, 75, 10]), but only provide annotations in the form of 2D bounding boxes, segmentation masks, or action labels. Some datasets for 3D hand pose estimation (e.g., [44, 76, 47]) include images of hands interacting with objects, but do not provide 6DoF object pose annotations.
@@
Similar to HOT3D, ground-truth poses of hands and objects in the recent ARCTIC dataset [15] were collected with a marker-based motion-capture system. This dataset includes 2.1M RGB images showing 10 subjects interacting with 11 articulated objects. The images were captured at 233K timestamps from 9 views, only one of which is egocentric (recorded with a mock-up of an egocentric device – a camera mounted on a helmet).
@@ 3 HOT3D dataset
833 minutes of recordings. The HOT3D dataset includes egocentric, multi-view, synchronized data streams recorded with Project Aria [13] and Quest 3 [41]. Image streams contain 1.5M multi-view frames consisting of 3.7M images. Each frame from Aria consists of one RGB 1408$\times$1408 image and two monochrome 640$\times$480 images. Each frame from Quest 3 consists of two monochrome 1280$\times$1024 images. Intrinsic camera parameters and camera-to-world transformations are available for all images. All streams were recorded at 30 fps. Every recording from Aria also includes a 3D point cloud of the scene (from SLAM) and per-frame eye gaze information. See supplement for more details about Aria and Quest 3.
@@
4 everyday scenarios. Besides a simple inspection scenario, where subjects pick up, observe, and put down objects, subjects were asked to perform typical actions in a kitchen, office, and living room. All scenarios were captured in the same lab equipped with scenario-specific furniture. In each recording, subjects were asked to interact with up to 6 objects. To enhance diversity within the dataset, we regularly randomized various aspects such as lighting, furniture placement, and decorative elements. The resulting dataset consists of 425 recordings, with 199 captured using Aria and 226 using Quest 3. Each recording is around 2 minutes long.
Ground-truth annotations. Recordings are annotated with per-frame ground-truth poses of hands and objects obtained in a motion-capture lab shown in Fig. 5 and described in the supplement. Object and wrist poses are represented as 3D rigid transformations from the 3D model space to the scene space, and hand poses are represented in the UmeTrack [26] and MANO [56] formats (UmeTrack is more accurate while MANO is more standard). Annotations in some frames may be missing or be of a lower quality. We visually inspected all recordings and flagged all frames with lower-quality poses. Out of 1.5M frames included in the dataset, 1.16M frames are fully annotated (i.e., ground-truth poses of all hands and objects are available) and passed our inspection. We release all 1.5M frames, which may be useful for unsupervised training, and provide a mask of the valid frames. See Fig. 4 and the supplement for statistics of the ground-truth object poses.
@@
Curated clips. To facilitate benchmarking of various tracking and pose estimation methods, we also release 3832 curated clips extracted from the full recordings. 2804 clips come from the training split and 1028 from the test split. Each clip has 150 frames (5 seconds) with ground-truth annotations available for all modeled objects and hands and passing our visual inspection.
Object onboarding sequences. To enable benchmarking model-free object tracking methods [62], which learn new objects from reference images, and 3D object reconstruction methods [42], HOT3D includes two types of onboarding sequences which show all possible views of each object: (1) sequences showing a static object on a desk, when the object is standing upright and upside-down, and (2) sequences showing an object manipulated by hands. The static onboarding setup is suitable for NeRF-like reconstruction methods [42], while the latter is more practical for AR/VR applications yet more challenging [22]. The ground-truth object poses are provided for all frames of the static sequences, but only for the first frame of the dynamic sequences. This is to simulate real-world settings, where the poses can be easily obtained by SfM [59] in the static setup, but are challenging to obtain in the dynamic setup. The ground-truth pose for the first frame of dynamic sequences is provided to define the canonical object space, which is necessary for evaluating 6DoF object tracking. Ground-truth hand poses are not provided for these sequences.
@@ 4 Experiments
Wearable headsets often feature multiple cameras, which naturally lends itself to the development of multi-view methods for 3D tasks. In this section, we demonstrate that multi-view methods outperform single-view methods for several popular egocentric tasks. First, we compare the single-view and multi-view versions of the UmeTrack [26] method for 3D hand tracking. Second, we extend the FoundPose [50] method for 6DoF pose estimation of unseen objects to multiple views and evaluate against the original version. Third, we develop a method for 3D lifting of unknown in-hand objects by stereo matching of DINOv2 [49] features, which we compare against a single-view method similar to OSNOM [51].
@@
Experimental setup. The task is to estimate 3D locations of hand joints in every frame of given test sequences, with the ground-truth hand skeletons and 2D bounding boxes of visible hands provided. We train the UmeTrack [26] hand tracker on three variants of training data: (1) training sequences from the UmeTrack dataset, (2) HOT3D training sequences recorded with Quest 3, and (3) the combination of the two. The UmeTrack dataset was recorded with the Quest 2 headset, which has the same but differently arranged cameras compared to Quest 3, and includes 1397 real and 1397 synthetic sequences, each recorded at 30 fps for 15 seconds. The sequences depict single-hand motions and hand-hand interactions performed by 53 participants, but do not include any hand-object interactions. All three UmeTrack models were trained on two-view image streams with one of the views randomly masked out (as in [26]). The masking increases the tracking robustness and encourages the models not to rely on both views, which enables a fair comparison of their single- and two-view modes. We evaluate the models on all frames of test UmeTrack sequences and all frames of test HOT3D clips from Quest 3. The accuracy of the predicted 3D locations is measured by the Mean Keypoint Position Error (MPKE) [26].
Results (Table 2). When trained only on the UmeTrack dataset, the UmeTrack tracker performs poorly on HOT3D, especially on frames with hand-object interactions that are not present in the UmeTrack dataset (24.2 MKPE on HOT3D vs. 13.6 on UmeTrack in the single-view mode). Conversely, we see a similar accuracy drop on the UmeTrack dataset when the tracker is trained on HOT3D, mostly on frames with hand-hand interactions that are not present in HOT3D (23.7 MKPE on UmeTrack vs. 18.0 on HOT3D). The accuracy drop is even larger when the tracker is trained only on one of the datasets and evaluated on two-view sequences. This is because the datasets are recorded with different headsets (Quest 2 vs. Quest 3) and the tracker overfitted to the camera configuration seen at training time. The domain gap between the two datasets is effectively closed when the tracker is trained on both, achieving 13.4 MKPE on UmeTrack and 15.4 on HOT3D in the single-view mode. A significant $41\%$ improvement (to 9.5 MKPE on UmeTrack and 10.9 on HOT3D) is then achieved when the model uses both views as input.
@@
Figure 6: Example 3D hand pose tracking results by UmeTrack [26]. Shown are hand skeletons and meshes of the UmeTrack hand model in the ground-truth and estimated poses. The UmeTrack hand tracker was provided participant’s ground-truth hand skeleton and was tasked to estimate 3D locations of the skeleton joints.
4.2 6DoF pose estimation of unseen objects
Experimental setup. In this experiment we focus on the task of model-based 6DoF pose estimation (i.e., 3D translation and 3D rotation) of unseen objects, which are learned during a short onboarding stage just from their CAD models [34]. We evaluate the original FoundPose [50], a recent open-source method that achieves the state-of-the-art results in refinement-free pose estimation from a single image, and its extension to multi-view input that we propose below. We evaluate coarse versions of these methods (without any pose refinement) on every 30th frame of the test HOT3D clips from both Aria and Quest 3. The methods are provided ground-truth 2D segmentation masks of visible objects and are tasked to estimate the 6DoF object poses. The accuracy of the estimated poses is measured by a recall rate defined as the fraction of samples for which a correct pose was estimated. A pose estimate is considered correct if its symmetry-aware translational and rotational errors are below a threshold.
@@
FoundPose [50] and its multi-view extension. During a short onboarding stage, FoundPose renders RGB-D templates showing the object model in different orientations, extracts DINOv2 patch features from the RGB channels, and registers the features in 3D using the depth channel. At inference time, FoundPose crops the RGB query image around the object mask and retrieves a small set of most similar templates using a DINOv2-based bag-of-words approach. For each retrieved template, a pose hypothesis is generated by PnP-RANSAC from 2D-3D correspondences established by matching patch features of the image crop and the template. Finally, the pose hypothesis with the highest number of inlier correspondences is selected as the pose estimate.
To study the effect of multi- vs. single-view input for this task, we propose a straightforward multi-view extension of the original FoundPose method. The extended version relies on the same onboarding stage, but at inference time crops the object in all available views, retrieves a handful of templates with the highest sum of per-crop cosine similarities, establishes multi-view 2D-3D correspondences between each of the templates and all views, and solves for the pose by solving the generalized PnP problem [37].
Results (Table 3). Our straightforward multi-view extension of FoundPose significantly outperforms the original single-view version, achieving 8–12% higher recall rates (13–34% relative improvement) on data from both headsets. Besides introducing additional constraints for 3D reasoning, the multi-view input offers more opportunities to observe objects that may be heavily or fully occluded in a single view. It is noteworthy that FoundPose performs well on the three-view setup from Aria, where one view is captured by a high-resolution RGB camera and the other two by lower-resolution monochrome cameras. We attribute this ability to generalize across different sensors primarily to DINOv2, which serves as the foundation model for FoundPose.
@@
Results (Table 4). The EgoHOS [75] model exhibits a noticeable decline in accuracy on the HOT3D dataset compared to its performance on the EgoHOS dataset, which we attribute to the domain gap between the training datasets (different camera configurations/models). This is particularly evident in the $\sim 50\%$ lower accuracy on Quest 3 monochrome frames. MRCNN-DA is the top performing method, outperforming EgoHOS on Aria and Quest 3 frames by 30% and 65% respectively. The incorporation of predicted depth maps enables more accurate disambiguation of foreground in-hand objects from the background, resulting in improved segmentation masks (Fig. 8).
@@
4.4 3D lifting of in-hand objects
Experimental setup. Next we focus on the task of 3D lifting of unknown in-hand objects, which is useful for object indexing and long-term tracking [51]. Given per-view 2D segmentation masks of an in-hand object, the goal is to estimate the 3D object location in the camera coordinate frame. We developed and compare the performance of three methods described below. The accuracy of the estimated locations is measured by a recall rate defined as the fraction of samples for which a correct location was estimated. A location is considered correct if its offset from the ground-truth location, defined by the center of the 3D object bounding box, is below a threshold.
Using hand as the object proxy (HandProxy). As a simple baseline, we use the ground-truth 3D palm center as the object location. We calculate the palm center as the middle point between the wrist joint and the first joint of the middle finger. If both hands are visible, we use the 3D palm location whose 2D projection is closer to the 2D object mask centroid. The goal of this baseline is to evaluate whether a specialized solution is necessary for this task or whether an accurate 3D hand tracker could provide a sufficient estimate.
Lifting by monocular depth estimation (MonoDepth). Our second approach is inspired by [51] and relies on monocular depth prediction and sparse SLAM observations. Specifically, we predict a monocular depth map by applying Depth Anything V2 [72] to a rectilinear input image (warped from the fisheye camera to a pinhole camera), align it to the 3D point cloud from SLAM by a scale-shift transformation [51], and finally calculate the 3D object location as the median of the aligned depth values collected from the object mask. Since MonoDepth requires SLAM observations that are not available for Quest 3, we evaluate this method only on recordings from Aria (using RGB images).

### v2

@@ Abstract
We introduce HOT3D, a publicly available dataset for egocentric hand and object tracking in 3D. The dataset offers over 833 minutes (3.7M+ images) of recordings that feature 19 subjects interacting with 33 diverse rigid objects. In addition to simple pick-up, observe, and put-down actions, the subjects perform actions typical for a kitchen, office, and living room environment. The recordings include multiple synchronized data streams containing egocentric multi-view RGB/monochrome images, eye gaze signal, scene point clouds, and 3D poses of cameras, hands, and objects. The dataset is recorded with two headsets from Meta: Project Aria, which is a research prototype of AI glasses, and Quest 3, a virtual-reality headset that has shipped millions of units. Ground-truth poses were obtained by a motion-capture system using small optical markers attached to hands and objects. Hand annotations are provided in the UmeTrack and MANO formats, and objects are represented by 3D meshes with PBR materials obtained by an in-house scanner. In our experiments, we demonstrate the effectiveness of multi-view egocentric data for three popular tasks: 3D hand tracking, model-based 6DoF object pose estimation, and 3D lifting of unknown in-hand objects. The evaluated multi-view methods, whose benchmarking is uniquely enabled by HOT3D, significantly outperform their single-view counterparts.
@@ 1 Introduction
Figure 2: Sample images from Aria (top) and Quest 3 (bottom). Aria recordings include one RGB and two monochrome image streams, while Quest 3 recordings include two monochrome streams (only images from one of the multi-view streams are shown). Contours of 3D models of hands and objects in the ground-truth poses are shown in white and green respectively. In addition to simple pick-up/observe/put-down actions, the subjects perform actions that are common in a kitchen, office, and living room. To increase diversity, the lighting, furniture, and decorations in the capture lab were regularly randomized.
@@
To accelerate computer vision research on hand-object interaction, we are publicly releasing HOT3D, an egocentric dataset recorded using two recent head-mounted devices from Meta: Project Aria [13], which is a research prototype of light-weight AI glasses, and Quest 3 [41], a virtual-reality headset that has shipped millions of units. The dataset offers over 833 minutes of egocentric video streams, including over 1.5M multi-view frames (3.7M+ images) and showing 19 subjects interacting with 33 diverse rigid objects. Besides a simple inspection scenario, where subjects pick up, observe, and put down the objects, the recordings show scenarios resembling typical actions in kitchen, office, and living room spaces. Hands and objects are annotated with accurate 3D poses collected using a passive marker-based motion-capture system. The dataset also includes 3D object models which were obtained by an in-house scanning-based 3D object reconstruction pipeline and include high-resolution geometry and PBR materials [40]. Recordings from Aria additionally include 3D scene point clouds from SLAM and eye gaze signal. Sample images are in Fig. 2.
The HOT3D dataset is primarily intended for the training and evaluation of hand and object tracking methods in 3D space from localized, egocentric, multi-view video streams, as opposed to monocular views or individual images. Since images from all streams are synchronized with a hardware trigger (i.e., captured at the same timestamp), the dataset enables development of methods that leverage multi-view and/or temporal information. While providing a testbed for CAD-based object tracking, the dataset also enables a CAD-free tracking setup by providing reference sequences showing different views at each object, which can be used to onboard the objects in a few-shot manner. Furthermore, the dataset can be used for tasks such as 3D object reconstruction and 2D detection or segmentation of hand-object interactions. We also encourage research that leverages the eye gaze information from Aria, which can enable predicting the user’s intent or efficient allocation of the computational budget via foveated sensing.
@@
HO-Cap [69] 699K RGB-D 1 / 9 HoloLens 9 Both 64 4 RGB-D + optim.
@@
Table 1: Datasets with 3D hands and object annotations. HOT3D is the first dataset to provide multi-view, hardware time-synced, egocentric videos captured with real headsets. It is the largest dataset in terms of total image count and provides high quality ground-truth annotations (comparable to [15]).
In our experiments, we primarily focus on demonstrating the effectiveness of multi-view egocentric data for several popular tasks. This type of data has been largely unexplored despite being commonly available on current AR/VR devices. Our results show that multi-view methods for 3D hand tracking, model-based 6DoF object pose estimation, and 3D lifting of unknown in-hand objects significantly outperform their single-view variants.
@@
    Strong baselines for tasks enabled by HOT3D: We developed simple yet powerful multi-view baseline methods for two tasks that are relevant for AR/VR and contextual AI applications: model-based 6DoF object pose estimation (extending FoundPose [50]), and 3D lifting of unknown in-hand objects (based on stereo matching of DINOv2 [49] features).
@@
    Demonstrated effectiveness of multi-view egocentric data: Our experiments show that multi-view methods for 3D hand tracking, 6DoF object pose estimation, and 3D lifting of handheld objects clearly outperform their single-view counterparts. This is an important result for research on power-efficient egocentric vision systems, which can typically afford a multi-view camera setup [13] but not, e.g., active depth sensors [68].
@@ 2 Related work
The progress of research in computer vision has been strongly influenced by benchmark datasets [58, 14, 57, 19, 32] which enable comparing methods and understanding their limitations. In this section, we first review existing datasets with either hand or object pose annotations, and then focus on datasets that offer annotations of hands and hand-manipulated objects.
Datasets with hands only. Vision-based 3D hand pose estimation and tracking has been studied for many years, with the first methods focusing on custom datasets with monochrome images [54, 29]. Significant improvements were later achieved on RGB-D images from datasets such as NYU [66], ICVL [65], MSRA [63], Tzionas et al. [67], EgoDexter [44], or HANDS17 [73]. Recently, partly motivated by AR/VR use cases where depth sensors are often unavailable due to high power consumption, the research community has largely switched to RGB or monochrome images, working on datasets such as Stereo [74], InterHand2.6M [43], FreiHAND [76], UmeTrack [26], AssemblyHands [47], and datasets with pose annotations of both hands and objects reviewed below.
@@
Datasets with hands and objects. Many existing datasets include images of hands and objects (e.g., [16, 6, 1, 60, 75, 10, 52]), but only provide annotations in the form of 2D bounding boxes, segmentation masks, or action labels. Some datasets for 3D hand pose estimation (e.g., [44, 76, 47]) include images of hands interacting with objects, but do not provide 6DoF object pose annotations.
@@
Similar to HOT3D, a marker-based motion-capture system was used to collect ground-truth poses of hands and objects in the recent ARCTIC dataset [15]. This dataset includes 2.1M RGB images showing 10 subjects interacting with 11 articulated objects. The images were captured at 233K timestamps from 9 views, only one of which is egocentric – recorded with a mock-up of an egocentric device (a camera mounted on a helmet).
Besides HOT3D, HO-Cap [69] is the only other dataset that is recorded with a real headset (Microsoft HoloLens) and provides 3D annotations of both hands and multiple objects, although the annotations were obtained using RGB-D cameras (instead of a precise motion-capture lab) and are therefore of a lower accuracy.
@@ 3 HOT3D dataset
833 minutes of recordings. HOT3D includes egocentric, multi-view, synchronized data streams recorded with Aria [13] and Quest 3 [41]. Image streams are recorded at 30 fps and contain 1.5M+ multi-view frames consisting of 3.7M+ images. Each Aria frame consists of one RGB 1408$\times$1408 image and two monochrome 640$\times$480 images. Each Quest 3 frame consists of two monochrome 1280$\times$1024 images. Intrinsic parameters and camera-to-world transformations are available for all images. Aria recordings also include 3D scene point clouds (from SLAM) and eye gaze signal. See appendix for details.
@@
4 everyday scenarios. In addition to a simple inspection scenario, where subjects pick up, observe, and put down objects, subjects were asked to perform typical actions in a kitchen, office, and living room. All scenarios were captured in the same lab equipped with scenario-specific furniture. In each recording, subjects were asked to interact with up to 6 objects. To enhance diversity within the dataset, we regularly randomized various aspects such as lighting, furniture placement, and decorative elements. The resulting dataset consists of 425 recordings, with 198 captured with Aria and 226 with Quest 3. Each recording is around 2 minutes long.
Ground-truth annotations. Recordings are annotated with per-frame ground-truth poses of hands and objects obtained in a motion-capture lab shown in Fig. 5 and described in the supplement. Object and wrist poses are represented as 3D rigid transformations from the 3D model space to the scene space, and hand poses are represented in the UmeTrack [26] and MANO [56] formats (UmeTrack is more accurate while MANO is more standard). Annotations in some frames may be missing or be of a lower quality. Out of 1.5M frames included in the dataset, 1.16M frames are fully annotated (i.e., ground-truth poses of all hands and objects are available) and passed our visual inspection (manually flagging frames were rendering of hand and object models in the ground-truth poses is not closely aligned with the observed image). We release all 1.5M frames, which may be useful for unsupervised training, and provide a mask of the valid 1.16M frames. See Fig. 4 and the appendix for statistics of the ground-truth object poses.
@@
Curated clips (HOT3D-Clips). To facilitate benchmarking of various tracking and pose estimation methods, we also release 3832 curated clips extracted from the full recordings (2804 clips come from the training and 1028 from the test split; 1983 from Aria and 1849 from Quest 3). Each clip has 150 frames (5 seconds) which all passed several quality-assurance tests (verifying that all hand and object annotations are present, at least one hand and one object are visible, discarding overexposed frames) and our visual inspection mentioned earlier.
Object onboarding sequences. To enable benchmarking model-free object tracking methods [62], which learn new objects from reference images, and 3D object reconstruction methods [42], HOT3D includes two types of onboarding sequences which show all possible views of each object: (1) sequences showing a static object on a desk, when the object is standing upright and upside-down, and (2) sequences showing an object manipulated by hands. The static onboarding setup is suitable for NeRF-like reconstruction methods [42], while the latter is more practical for AR/VR applications yet more challenging [22]. The ground-truth object poses are provided for all frames of the static sequences, but only for the first frame of the dynamic sequences. This is to simulate real-world settings, where the poses can be easily obtained by SfM [59] in the static setup, but are challenging to obtain in the dynamic setup. The ground-truth pose for the first frame of dynamic sequences is provided to define the canonical object space, which is necessary for evaluating 6DoF object tracking.
@@ 4 Experiments
Wearable headsets often feature multiple cameras, which makes them naturally suited for developing multi-view 3D vision methods. In this section, we demonstrate that multi-view methods outperform single-view methods for several popular egocentric tasks. First, we compare the single-view and multi-view versions of the UmeTrack [26] method for 3D hand tracking. Second, we extend the FoundPose [50] method for model-based 6DoF object pose estimation to multiple views and evaluate against the original version. Third, we develop a method for 3D lifting of unknown in-hand objects by stereo matching of DINOv2 [49] features, which we compare against a single-view method similar to OSNOM [51]. Note that all 3D predictions in our experiments were expressed in the headset space but could be transformed to the world space as the camera-to-world transformation is available (from on-headset SLAM).
@@
Experimental setup. Given the hand shape (3D hand skeleton in the canonical pose) and ground-truth 2D bounding boxes of visible hands, the task is to estimate the hand poses (3D locations of skeleton joints) in every frame of an input sequence. We train the UmeTrack [26] hand tracker on three variants of training data: (1) training sequences from the UmeTrack dataset, (2) HOT3D training sequences recorded with Quest 3, and (3) the combination of the two. The UmeTrack dataset was recorded with the Quest 2 headset, which has the same but differently arranged cameras compared to Quest 3, and includes 1397 real and 1397 synthetic sequences, each recorded at 30 fps for 15 seconds. The sequences depict single-hand motions and hand-hand interactions performed by 53 participants, but do not include any hand-object interactions. All three UmeTrack models were trained on two-view image streams with one of the views randomly masked out (as in [26]). The masking increases the tracking robustness and encourages the models not to rely on both views, which enables a fair comparison of their single- and two-view modes. We evaluate the models on all frames of test UmeTrack sequences and all frames of test HOT3D clips from Quest 3. The accuracy of the predicted 3D locations is measured by the Mean Keypoint Position Error (MPKE) [26].
Results (Tab. 2). When trained on the UmeTrack dataset, the hand tracker in the single-view mode performs poorly on HOT3D (24.2 MKPE on HOT3D vs. 13.6 on UmeTrack). Similarly, when trained on HOT3D, the single-view tracker performs poorly on UmeTrack (23.7 MKPE on UmeTrack vs. 18.0 on HOT3D). The main reason of these accuracy gaps is that hand-object interactions are present only in HOT3D while hand-hand interactions only in the UmeTrack dataset. The accuracy drop is even larger when the tracker, still trained only on one of the datasets, is evaluated in the two-view mode. This is because the datasets are recorded with different headsets (Quest 2 vs. Quest 3) and the tracker overfits to the camera configuration seen at training. The domain gap between the two datasets is effectively closed when the tracker is trained on both, achieving 13.4 MKPE on UmeTrack and 15.4 on HOT3D in the single-view mode, and a significant $41\%$ improvement (9.5 MKPE on UmeTrack and 10.9 on HOT3D) in the two-view mode.
@@
Figure 6: Example 3D hand pose tracking results by UmeTrack [26]. Shown are hand skeletons and meshes of the UmeTrack hand model in the ground-truth and estimated poses. The UmeTrack hand tracker was provided the ground-truth hand skeleton and was tasked to estimate 3D locations of the skeleton joints.
4.2 6DoF object pose estimation
Experimental setup. In this experiment we focus on training-free 6DoF object pose estimation, where objects are onboarded during a short stage using only their CAD models [34]. We evaluate FoundPose [50], a recent open-source method that achieves state-of-the-art results in refinement-free pose estimation from a single RGB image, and its extension to multi-view input that we propose below. We evaluate refinement-free versions of these methods on every 30th frame of the test HOT3D clips from both Aria and Quest 3. The input of the methods is a single/multi-view frame with ground-truth 2D segmentation masks of visible objects. The accuracy of the estimated poses is measured by a recall rate defined as the fraction of samples for which a correct pose was estimated. A pose estimate is considered correct if its symmetry-aware translational and rotational errors are below a threshold.
@@
FoundPose [50] and its multi-view extension. During a short onboarding stage, FoundPose renders RGB-D templates showing the object model in different orientations, extracts DINOv2 patch features from the RGB channels, and registers the features in 3D using the depth channel. At inference time, FoundPose crops the RGB query image around the object mask and retrieves a small set of most similar templates using a DINOv2-based bag-of-words approach. For each retrieved template, a pose hypothesis is generated by PnP-RANSAC from 2D-3D correspondences established by matching DINOv2 patch features of the image crop and the template. Finally, the pose hypothesis with the highest number of inlier correspondences is selected as the pose estimate.
To study the effect of multi- vs. single-view input for this task, we propose a straightforward multi-view extension of the original FoundPose method. The extended version relies on the same onboarding stage, but at inference time crops the object in all available views, retrieves a handful of templates with the highest sum of per-view template scores, establishes multi-view 2D-3D correspondences between each of the templates and all views, and solves for the pose by solving the generalized PnP problem [37].
Results (Tab. 3). Our straightforward multi-view extension of FoundPose significantly outperforms the original single-view version, achieving 8–12% higher recall rates (13–34% relative improvement) on data from both headsets. Besides introducing additional constraints for 3D reasoning, the multi-view input offers more opportunities to observe objects that may be heavily or fully occluded in a single view. It is noteworthy that FoundPose performs well on the two-view monochrome setup from Quest3 and also the three-view setup from Aria, where one view is captured by a high-resolution RGB camera and the other two by lower-resolution monochrome cameras. We attribute this ability to generalize across different sensors primarily to DINOv2, which serves as a backbone for FoundPose.
@@
Results (Tab. 4). The EgoHOS [75] model exhibits a noticeable decline in accuracy on HOT3D compared to its performance on the EgoHOS dataset, which we attribute to the domain gap between the datasets. This is particularly evident in the $\sim 50\%$ lower accuracy on Quest 3 monochrome images. MRCNN-DA is the top performing method, outperforming EgoHOS on Aria and Quest 3 frames by 30% and 65% respectively. The incorporation of predicted depth maps enables more accurate disambiguation of foreground in-hand objects from the background, resulting in improved segmentation masks (Fig. 8).